# and Aggregation Improves Adversarial Robustness

### v5

@@ Abstract
In this work, we propose a novel adversarial defence mechanism for image classification – Carso – blending the paradigms of adversarial training and adversarial purification in a synergistic robustness-enhancing way. The method builds upon an adversarially-trained classifier, and learns to map its internal representation associated with a potentially perturbed input onto a distribution of tentative clean reconstructions. Multiple samples from such distribution are classified by the same adversarially-trained model, and a carefully chosen aggregation of its outputs finally constitutes the robust prediction of interest. Experimental evaluation by a well-established benchmark of strong adaptive attacks, across different image datasets, shows that Carso is able to defend itself against adaptive end-to-end white-box attacks devised for stochastic defences. Paying a modest clean accuracy toll, our method improves by a significant margin the state-of-the-art for Cifar-10, Cifar-100, and TinyImageNet-200 $\ell_{\infty}$ robust classification accuracy against AutoAttack.
@@ 1 Introduction
Given the widespread nature of the issue (Ilyas et al., 2019), and the serious concerns raised about the safety and reliability of data-learnt models in the lack of an appropriate mitigation (Biggio & Roli, 2018), adversarial attacks have been extensively studied. However, obtaining generally robust machine learning (ML) systems remains a longstanding issue, and a major open challenge.
@@
Research in the field has been driven by two opposing, yet complementary, efforts. On the one hand, the study of failure modes in existing models and defences, with the goal of understanding their origin and developing stronger attacks with varying degrees of knowledge and control over the target system (Szegedy et al., 2014; Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016; Tramèr et al., 2020). On the other hand, the construction of increasingly capable defence mechanisms. Although alternatives have been explored (Cisse et al., 2017; Tramèr et al., 2018; Carbone et al., 2020; Zhang et al., 2022), most of the latter is based on adequately leveraging adversarial training (Goodfellow et al., 2015; Madry et al., 2018; Tramèr & Boneh, 2019; Rebuffi et al., 2021; Gowal et al., 2021; Jia et al., 2022; Singh et al., 2023; Wang et al., 2023; Cui et al., 2023; Peng et al., 2023), i.e. training a ML model on a dataset composed of (or enriched with) adversarially-perturbed inputs associated with their correct, pre-perturbation labels. In fact, adversarial training has been the only technique capable of consistently providing an acceptable level of defence (Gowal et al., 2020), while still incrementally improving up to the current state-of-the-art (Cui et al., 2023; Peng et al., 2023; Bartoldson et al., 2024).
@@
Another defensive approach is that of adversarial purification (Shi et al., 2021; Yoon et al., 2021), where a generative model is used – similarly to denoising – to recover a perturbation-free version of the input before classification is performed. Nonetheless, such attempts have generally fallen short of expectations due to inherent limitations of the generative models used in early attempts (Nie et al., 2022), or due to decreases in robust accuracy [^1] The test set accuracy of the frozen-weights trained classifier – computed on a dataset entirely composed of adversarially-perturbed examples generated against that specific model. when attacked end-to-end (Gu & Rigazio, 2015) – resulting in subpar robustness if the defensive structure is known to the adversary (Tramèr et al., 2020). More recently, the rise of diffusion-based generative models (Huang et al., 2021) and their use for purification have enabled more successful results of this kind (Nie et al., 2022; Chen et al., 2023) – although at the cost of longer training and inference times, and a much brittler robustness evaluation (Chen et al., 2023; Lee & Kim, 2024).
@@
In this work, we design a novel adversarial defence for supervised image classification, dubbed Carso (i.e., Counter-Adversarial Recall of Synthetic Observations). The approach relies on an adversarially-trained classifier (called hereinafter simply the classifier), endowed with a stochastic generative model (called hereinafter the purifier). Upon classification of a potentially-perturbed input, the latter learns to generate – from the tensor [^2] Which we call internal representation. of (pre)activations registered at neuron level in the former – samples from a distribution of plausible, perturbation-free reconstructions. At inference time, some of these samples are classified by the very same classifier, and the original input is robustly labelled by aggregating its many outputs in the form of a normalised doubly-exponential logit product. This method – to the best of our knowledge the first attempt to organically merge the adversarial training and purification paradigms – avoids the vulnerability pitfalls typical of the mere stacking of a purifier and a classifier (Gu & Rigazio, 2015), while still being able to take advantage of independent incremental improvements to adversarial training or generative modelling.
@@
An empirical assessment [^3] Implementation of the method and code for the experiments (based on PyTorch (Paszke et al., 2019), AdverTorch (Ding et al., 2019), and ebtorch (Ballarin, 2024)) can be found at: https://github.com/emaballarin/CARSO. of the defence in the $\ell_{\infty}$ white-box setting is provided, using a conditional (Sohn et al., 2015; Yan et al., 2016) variational autoencoder (Kingma & Welling, 2014; Rezende et al., 2014) as the purifier and existing state-of-the-art adversarially pre-trained models as classifiers. Such choices are meant to give existing approaches – and the adversary attacking our architecture end-to-end as part of the assessment – the strongest advantage possible. Yet, in all scenarios considered, Carso improves significantly the robustness of the pre-trained classifier – even against attacks specifically devised to fool stochastic defences like ours. Remarkably, with a modest clean accuracy toll, our method improves by a significant margin the current state-of-the-art for Cifar-10 (Krizhevsky, 2009), Cifar-100 (Krizhevsky, 2009), and TinyImageNet-200 (Chrabaszcz et al., 2017) $\ell_{\infty}$ robust classification accuracy against AutoAttack (Croce & Hein, 2020a).
@@
The rest of the manuscript is structured as follows. In section 2 we provide an overview of selected contributions in the fields of adversarial training and purification-based defences – with focus on image classification. In section 3, a deeper analysis is given of two integral parts of our experimental assessment: Pgd adversarial training and conditional variational autoencoders. Section 4 is devoted to the intuition behind Carso, its architectural description, and the relevant technical details that allow it to work effectively. Section 5 contains details about the experimental setup, results, comments, and limitations. Section 6 concludes the paper and outlines directions of future development.
@@ 2 Related work
The introduction of Fgsm (Goodfellow et al., 2015) enables the efficient generation of adversarial examples along the training, with a single normalised gradient step. Its iterative counterpart Pgd (Madry et al., 2018) – discussed in section 3 and Appendix A – significantly improves the effectiveness of the adversarial examples produced, making it still the de facto standard for the synthesis of adversarial training inputs (Gowal et al., 2020). Further incremental improvements have also been developed, some focused specifically on robustness assessment (e.g. stepsize-adaptive variants, as by Croce & Hein (2020a)).
@@ 3 Preliminaries
The formal details of such method are provided in Appendix A.
@@
Variational autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) allow the learning from data of approximate generative latent-variable models of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, whose likelihood and posterior are approximately parameterised by deep artificial neural networks (ANNs). The problem is cast as the maximisation of a variational lower bound.
In practice, optimisation is performed iteratively – on a loss function given by the linear mixture of data-reconstruction loss and empirical KL divergence w.r.t. a chosen prior, computed on mini-batches of data.
@@
Further details on the functioning of such models are given in Appendix B.
@@ 4.1 Architectural overview and principle of operation
There are no specific requirements for the classifier, whose training is completely independent of the use of the model as part of Carso. However, training it adversarially significantly improves the robust accuracy of the overall system, also allowing it to benefit from established adversarial training techniques.
The purifier is also independent of specific architectural choices, provided it is capable of stochastic conditional data generation at inference time, with the internal representation of the classifier used as the conditioning set.
@@
In fact, the decoder of a conditional VAE allows for exact algorithmic differentiability (Baydin et al., 2018) w.r.t. its conditioning set, thus averting the need for backward-pass approximation (Athalye et al., 2018a) in generating end-to-end adversarial attacks against the entire system, and preventing (un)intentional robustness by gradient obfuscation (Athalye et al., 2018a). The same cannot be said (Chen et al., 2023) for more capable and modern purification models, such as those based e.g. on diffusive processes, whose robustness assessment is still in the process of being thoroughly understood (Lee & Kim, 2024).
A downside of such choice is represented by the reduced effectiveness of the decoder in the synthesis of complex data, due to well-known model limitations. In fact, we experimentally observe a modest increase in reconstruction cost for non-perturbed inputs, which in turn may limit the clean accuracy of the entire system. Nevertheless, we defend the need for a fair and transparent robustness evaluation, such as the one provided by the use of a VAE-based purifier, in the evaluation of any novel architecture-agnostic adversarial defence technique.
A diagram of the whole architecture is shown in Figure 1, and its detailed principles of operation are recapped below.
Figure 1: Schematic representation of the Carso architecture used in the experimental phase of this work. The subnetwork bordered by the red dashed line is used only during the training of the purifier. The subnetwork bordered by the blue dashed line is re-evaluated on different random samples $\bm{z}_{i}$ and the resulting individual $\hat{y}_{i}$ are aggregated into $\hat{y}_{\text{rob}}$. The classifier $f(\cdot;\bm{\theta})$ is always kept frozen; the remaining network is trained on $\mathcal{L}_{\text{VAE}}(\bm{x},\hat{\bm{x}})$. More precise details on the functioning of the networks are provided in subsection 4.1.
@@
At training time, adversarially-perturbed examples are generated against the classifier, and fed to it. The tensors containing the classifier (pre)activations across the network are then extracted. Finally, the conditional VAE serving as the purifier is trained on perturbation-free input reconstruction, conditional on the corresponding previously extracted internal representations, and using pre-perturbation examples as targets.
@@ 4.3 Hierarchical input and internal representation encoding
In order to use the VAE decoder at inference time, the entire compression machinery for the conditioning set must be preserved after training, and used to encode the internal representations extracted. The input encoder may be discarded instead.
@@ 4.4 Adversarially-balanced batches
Any smaller subset of attack types and strengths, or a detailedly unbalanced batch composition, always experimentally results in a worse performing purification model. More details justifying such choice are provided in Appendix C.
@@ 4.5 Robust aggregation strategy
Such choice produces a robust prediction much harder to overtake in the event that an adversary selectively targets a specific input reconstruction. A heuristic analysis of this property (subsection D.1), together with an experimental justification for using such aggregation function within Carso (subsection D.2), are given in Appendix D.
@@ 5.1 Setup
A WideResNet-28-10 model is used as the classifier, adversarially pre-trained on the respective dataset – the only difference between scenarios being the number of output logits: $10$ in scenario (a), $100$ in scenario (b), and $200$ in scenario (c).
@@
More precise details on all architectures are given in Appendix E.
@@
The purifier is trained on the VAE loss, using summed pixel-wise channel-wise binary cross-entropy as the reconstruction cost. Optimisation is performed by RAdam+Lookahead (Liu et al., 2020; Zhang et al., 2019b) with a learning rate schedule that presents a linear warm-up, a plateau phase, and a linear annealing (Smith, 2017). To promote the learning of meaningful reconstructions during the initial phases of training, the KL divergence term in the VAE loss is suppressed for an initial number of epochs. Afterwards, it is linearly modulated up to its actual value, during a fixed number of epochs ($\beta$ increase) (Higgins et al., 2017). The initial and final epochs of such modulation are reported in Table 15.
Additional scenario-specific details are provided in Appendix E.
@@
The complete details and hyperparameters of the attacks are described in Appendix E.
@@
In each scenario, we report the clean and robust test-set accuracy – the latter by means of AutoAttack (Croce & Hein, 2020a) – of the classifier and the corresponding Carso architecture.
@@
In the evaluation of the Carso architecture, the number of reconstructed samples per input is set to $8$, the logits are aggregated as explained in subsection 4.5, and the output class is finally selected as the $\operatorname*{\arg\max}$ of the aggregation. Due to the stochastic nature of the purifier, robust accuracy is assessed by a version of AutoAttack suitable for stochastic defences (randAA) – composed of AutoPgd on the cross-entropy and difference of logits ratio losses, across $20$ Expectation over Transformation (EoT) (Athalye et al., 2018b) iterations with $100$ gradient ascent steps each. In the specific case of scenario (a), we also assess our method by the Pgd+EoT pipeline proposed by Lin et al. (2024), as explained in subsection 5.2.
@@ 5.2 Results and discussion
An analysis of the experimental results is provided in the subsection that follows, whereas their systematic exposition is given in Table 2.
Table 2: Clean (results in italic) and adversarial (results in upright) accuracy for the different models and datasets used in the respective scenarios. The following abbreviations are used: Scen: scenario considered; AT/Cl: clean accuracy for the adversarially-pretrained model used as the classifier, when considered alone; C/Cl: clean accuracy for the Carso architecture; AT/AA: robust accuracy (by the means of AutoAttack) for the adversarially-pretrained model used as the classifier, when considered alone; C/randAA: robust accuracy for the Carso architecture, when attacked end-to-end by AutoAttack for randomised defences; Best AT/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial training alone (any model); Best P/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial purification (any model). Robust accuracies in round brackets are obtained using the Pgd+EoT (Lee & Kim, 2024) pipeline, developed for diffusion-based purifiers. The best clean and robust accuracies per dataset are shown in bold. The clean accuracies for the models referred to in the Best columns are shown in Table 16 (in Appendix F).
@@
Comparing the robust accuracy of the classifier model used in scenario (a) (Cui et al., 2023) with that resulting from the inclusion of the same model in the Carso architecture, we observe a $+8.4\%$ increase. This is counterbalanced by a $-5.6\%$ clean accuracy toll. The same version of Carso further provides a $+2.42$ robustness increase w.r.t. the current best AT-trained model (Bartoldson et al., 2024) that employs a 4$\sim$ 4$\times$ 4 × larger WideResNet-96-16 model.
In addition, our method provides a remarkable $+9.72\%$ increase in robust accuracy w.r.t. to the best adversarial purification approach (Lin et al., 2024), a diffusion-based purifier. However, the comparison is not as straightforward. In fact, the original paper (Lin et al., 2024) reports a robust accuracy of $78.12\%$ using AutoAttack on the gradients obtained via the adjoint method (Nie et al., 2022). As noted in Lee & Kim (2024), such evaluation (which uses the version of AutoAttack that is unsuitable for stochastic defences) leads to a large overestimation of the robustness of diffusive purifiers. As suggested in Lee & Kim (2024), the authors of Lin et al. (2024) re-evaluate the robust accuracy according to a more suitable pipeline (Pgd+EoT, whose hyperparameters are shown in Table 13), obtaining a much lower robust accuracy of $66.41\%$. Consequently, we repeat the same evaluation for Carso and compare the worst-case robustness amongst the two. In line with typical AT methods, and unlike diffusive purification, the robustness of Carso assessed by means of randAA remains lower w.r.t. that achieved by Pgd+EoT.
@@
Moving to scenario (b), Carso achieves a robust accuracy increase of $+27.47\%$ w.r.t. the classifier alone (Cui et al., 2023), balanced by a $-5.79\%$ decrease in clean accuracy. Our approach also improves upon the robust accuracy of the best AT-trained model (Wang et al., 2023) (WideResNet-70-16) by $+23.98\%$. In the absence of a reliable robustness evaluation by means of Pgd+EoT for the best purification-based method (Lin et al., 2024), we still obtain a $+20.25\%$ increase in robust accuracy upon its (largely overestimated) AA result.
@@
In scenario (c), Carso improves upon the classifier alone (Wang et al., 2023) (which is also the best AT-based approach for TinyImageNet-200) by $+22.26\%$. A significant clean accuracy toll is imposed by the relative complexity of the dataset, i.e. $-8.87\%$. In this setting, we lack any additional purification-based methods.
@@
Although the architecture of Carso is algorithmically differentiable end-to-end – and the integrated diagnostics of the randAA routines raised no warnings during the assessment – we additionally guard against the eventual gradient obfuscation (Athalye et al., 2018a) induced by our method by repeating the evaluation at $\epsilon_{\infty}=0.95$, verifying that the resulting robust accuracy stays below random chance (Carlini et al., 2019). Results are shown in Table 3.
@@
Scenario (a) (b) (c) $\epsilon_{\infty}=0.95$ acc. $<$0.047 $<$0.010 $\approx$0.0

### v6

# Blending adversarial training and representation-conditional purification via aggregation improves adversarial robustness

@@ Abstract
In this work, we propose a novel adversarial defence mechanism for image classification – Carso – blending the paradigms of adversarial training and adversarial purification in a synergistic robustness-enhancing way. The method builds upon an adversarially-trained classifier, and learns to map its internal representation associated with a potentially perturbed input onto a distribution of tentative clean reconstructions. Multiple samples from such distribution are classified by the same adversarially-trained model, and a carefully chosen aggregation of its outputs finally constitutes the robust prediction of interest. Experimental evaluation by a well-established benchmark of strong adaptive attacks, across different image datasets, shows that Carso is able to defend itself against adaptive end-to-end white-box attacks devised for stochastic defences. With a modest clean accuracy penalty, our method improves by a significant margin the state-of-the-art for Cifar-10, Cifar-100, and TinyImageNet-200 $\ell_{\infty}$ robust classification accuracy against AutoAttack.
@@ 1 Introduction
Given the widespread nature of the issue (Ilyas et al., 2019), and the serious concerns raised about the safety and reliability of models learnt from data in the lack of appropriate mitigations (Biggio & Roli, 2018), adversarial attacks have been extensively studied. However, obtaining generally robust machine learning (ML) systems remains a longstanding issue, and a major open challenge.
@@
Research in the field has been driven by two opposing yet complementary efforts. On the one hand, the study of failure modes in existing models and defences, with the goal of understanding their origin and developing stronger attacks with varying degrees of knowledge and control over the target system (Szegedy et al., 2014; Goodfellow et al., 2015; Moosavi-Dezfooli et al., 2016; Tramèr et al., 2020). On the other hand, the construction of increasingly capable defence mechanisms. Although alternatives have been explored (Cisse et al., 2017; Tramèr et al., 2018; Carbone et al., 2020; Zhang et al., 2022), most of the latter is based on adequately leveraging adversarial training (Goodfellow et al., 2015; Madry et al., 2018; Tramèr & Boneh, 2019; Rebuffi et al., 2021; Gowal et al., 2021; Jia et al., 2022; Singh et al., 2023; Wang et al., 2023; Cui et al., 2023; Peng et al., 2023), i.e. training a ML model on a dataset composed of (or enriched with) adversarially-perturbed inputs associated with their correct, pre-perturbation labels. In fact, adversarial training has been the only technique capable of consistently providing an acceptable level of defence (Gowal et al., 2020), while still incrementally improving up to the current state-of-the-art (Cui et al., 2023; Peng et al., 2023; Bartoldson et al., 2024).
@@
Another defensive approach is that of adversarial purification (Shi et al., 2021; Yoon et al., 2021), where a generative model is used – similarly to denoising – to recover a perturbation-free version of the input before classification is performed. Nonetheless, such attempts have generally fallen short of expectations due to inherent limitations of the generative models used in early attempts (Nie et al., 2022), or due to decreases in robust accuracy [^1] The test set accuracy of the frozen-weights trained classifier – computed on a dataset entirely composed of adversarially-perturbed examples generated against that specific model. when attacked end-to-end (Gu & Rigazio, 2015; Lee & Kim, 2024) – resulting in subpar robustness if the defensive structure is known to the adversary (Tramèr et al., 2020). More recently, the rise of diffusion-based generative models (Huang et al., 2021) and their use for purification have enabled more successful results of this kind (Nie et al., 2022; Chen et al., 2023) – although at the cost of longer training and inference times, and a much brittler robustness evaluation (Chen et al., 2023; Lee & Kim, 2024).
@@
In this work, we design a novel adversarial defence for supervised image classification, dubbed Carso (Counter-Adversarial Recall of Synthetic Observations). The approach relies on an adversarially-trained classifier (called hereinafter simply the classifier), endowed with a stochastic generative model (called hereinafter the purifier). Upon classification of a potentially-perturbed input, the latter learns to generate – from the tensor [^2] Which we call internal representation. of (pre)activations registered at neuron level in the former – samples from a distribution of plausible, perturbation-free reconstructions. At inference time, some of these samples are classified by the very same classifier, and the original input is robustly labelled by aggregating its many outputs in the form of a normalised doubly-exponential logit product. This method – to the best of our knowledge the first attempt to organically merge the adversarial training and purification paradigms – avoids the vulnerability pitfalls typical of the mere stacking of a purifier and a classifier (Gu & Rigazio, 2015; Lee & Kim, 2024), while still being able to take advantage of independent incremental improvements to adversarial training or generative modelling.
@@
An empirical assessment [^3] Implementation of the method and code for the experiments (based on PyTorch (Paszke et al., 2019), AdverTorch (Ding et al., 2019), and ebtorch (Ballarin, 2025)) can be found at: https://github.com/emaballarin/CARSO. of the defence in the $\ell_{\infty}$ white-box setting is provided, using a conditional (Sohn et al., 2015; Yan et al., 2016) variational autoencoder (Kingma & Welling, 2014; Rezende et al., 2014) as the purifier and existing state-of-the-art adversarially pre-trained models as classifiers. Such choices are meant to give existing approaches – and the adversary attacking our architecture end-to-end as part of the assessment – the strongest advantage possible. Yet, in all scenarios considered, Carso improves significantly the robustness of the pre-trained classifier – even against attacks specifically devised to fool stochastic defences like ours. Remarkably, with a modest clean accuracy penalty, our method improves by a significant margin the current state-of-the-art for Cifar-10 (Krizhevsky, 2009), Cifar-100 (Krizhevsky, 2009), and TinyImageNet-200 (Chrabaszcz et al., 2017) $\ell_{\infty}$ robust classification accuracy against AutoAttack (Croce & Hein, 2020a).
@@
The rest of the manuscript is structured as follows. In section 2 we provide an overview of selected contributions in the fields of adversarial training and purification-based defences – with focus on image classification. In section 3, an introduction is given to two integral parts of our experimental assessment: Pgd adversarial training and conditional variational autoencoders. Section 4 is devoted to the intuition behind Carso, its architectural description, and the relevant technical details that allow it to work effectively. Section 5 contains details about the experimental setup, results, comments, and limitations. Section 6 concludes the paper and outlines directions of future development.
@@ 2 Related work
The introduction of Fgsm (Goodfellow et al., 2015) enables the efficient generation of adversarial examples along the training, with a single normalised gradient step. Its iterative generalisation Pgd (Madry et al., 2018) – discussed in section 3 – significantly improves the effectiveness of the adversarial examples produced, making it still the de facto standard for the synthesis of adversarial training inputs (Gowal et al., 2020). Further incremental improvements have also been developed, some focused specifically on robustness assessment (e.g. stepsize-adaptive variants, as by Croce & Hein (2020a)).
@@ 3 Preliminaries
Variational autoencoders (VAEs) (Kingma & Welling, 2014; Rezende et al., 2014) allow the learning from data of approximate generative latent-variable models of the form $p(\bm{x},\bm{z})=p(\bm{x}\operatorname*{|}\bm{z})p(\bm{z})$, whose likelihood and posterior are approximately parametrised by deep artificial neural networks (ANNs). The problem is cast as the maximisation of a variational lower bound.
In practice, optimisation is performed iteratively – on a loss function ($\mathcal{L}_{\text{VAE}}$) given by the linear mixture of data-reconstruction loss and empirical KL divergence w.r.t. a chosen prior, computed on mini-batches of data.
@@ 4.1 Architectural overview and principle of operation
There are no specific requirements for the classifier, whose training is completely independent of the use of the model as part of Carso. However, training it adversarially significantly improves the robust accuracy of the overall system (see Appendix D), also allowing it to benefit from established adversarial training techniques.
The purifier is also independent of specific architectural choices, provided it is capable of stochastic conditional data generation at inference time, with the internal representation of the classifier used as conditioning.
@@
In fact, the decoder of a conditional VAE allows for exact algorithmic differentiability (Baydin et al., 2018) w.r.t. its conditioning set, thus averting the need for backward-pass approximation (Athalye et al., 2018a) in generating end-to-end adversarial attacks against the entire system, and preventing (un)intentional robustness inflation by gradient obfuscation (Athalye et al., 2018a). The same cannot be said (Chen et al., 2023) for more capable and modern purification models, such as those based e.g. on diffusive processes, whose proper robustness assessment is still in the process of being thoroughly understood (Lee & Kim, 2024).
A downside of such choice is represented by the reduced effectiveness of the decoder in the synthesis of complex data, due to well-known model limitations. In fact, we experimentally observe a modest increase in reconstruction cost for non-perturbed inputs, which in turn may limit the clean accuracy of the entire system. Nevertheless, we defend the need for a fair and transparent robustness evaluation, such as the one provided by the use of a VAE-based purifier, in the evaluation of any novel architecture-agnostic adversarial defence.
A diagram of the whole architecture is shown in Figure 1, and its detailed principles of operation are recapped below. Additionally, an ablation study investigating the need for either the classifier or the purifier being trained on adversarially-perturbed inputs is provided in Appendix D.
[…]
@@
At training time, adversarially-perturbed examples are generated against the classifier, and fed to it. The tensors containing the classifier (pre)activations across the network are then extracted. Finally, the conditional VAE serving as the purifier is trained on perturbation-free input reconstruction, conditional on the corresponding previously-extracted internal representations, and using pre-perturbation examples as targets.
@@ 4.3 Hierarchical input and internal representation encoding
In order to use the VAE decoder at inference time, the compression machinery for the conditioning set must be preserved after training, and used to encode the internal representations extracted. The input encoder may be discarded instead.
@@ 4.4 Adversarially-balanced batches
Any smaller subset of attack types and strengths, or a detailedly unbalanced batch composition, experimentally results in a worse-performing purification model. More details justifying such choice are provided in Appendix A.
@@ 4.5 Robust aggregation strategy
Such choice produces a robust prediction much harder to overtake in the event that an adversary selectively targets a specific input reconstruction. A heuristic analysis of this property (subsection B.1), together with an experimental justification for using such aggregation function within Carso (subsection B.2), are given in Appendix B.
@@ 5.1 Setup
A WideResNet-28-10 model is used as the classifier, adversarially pre-trained on the respective dataset – the only difference between scenarios being the size of the inputs, and the number of output logits: $10$ in scenario (a), $100$ in scenario (b), and $200$ in scenario (c).
@@
More precise details on all architectures are given in Appendix C.
@@
The purifier is trained on the VAE loss, using summed pixel-wise channel-wise binary cross-entropy as the reconstruction cost. Optimisation is performed by RAdam+Lookahead (Liu et al., 2020; Zhang et al., 2019b) with a learning rate schedule that presents a linear warm-up, a plateau phase, and a linear annealing (Smith, 2017). To promote the learning of meaningful reconstructions during the initial phases of training, the KL divergence term in the VAE loss is suppressed for an initial number of epochs. Afterwards, it is linearly modulated up to its actual value, along a fixed number of epochs ($\beta$ increase) (Higgins et al., 2017). The initial and final epochs of such modulation are reported in Table 16.
Additional scenario-specific details are provided in Appendix C.
@@
The complete details and hyperparameters of the attacks are described in Appendix C.
@@
In each scenario, we report the clean and robust test-set accuracy – the latter by means of AutoAttack (Croce & Hein, 2020a) – of the classifier alone, and that of the corresponding Carso architecture.
@@
In the evaluation of the Carso architecture, the number of reconstructed samples per input is set to $8$, the logits are aggregated as explained in subsection 4.5, and the output class is finally selected as the $\operatorname*{\arg\max}$ of the aggregation. Due to the stochastic nature of the purifier, robust accuracy is assessed by a version of AutoAttack suitable for stochastic defences (randAA) – composed of AutoPgd on the cross-entropy and difference of logits ratio losses, across $20$ Expectation over Transformation (EoT) (Athalye et al., 2018b) iterations with $100$ gradient ascent steps each. In the specific case of scenario (a), we also assess our method by the Pgd+EoT pipeline proposed by Lee & Kim (2024), as explained in subsection 5.2.
@@ 5.2 Results and discussion
An analysis of the experimental results is provided in the subsection that follows, whereas their systematic exposition is given in Table 2. Results obtained by using deliberately worse-performing pretrained classifiers, as well as a broader comparison with existing adversarial defences from literature, are provided in Appendix E.
Table 2: Clean (results in italic) and adversarial (results in upright) accuracy for the different models and datasets used in the respective scenarios. The following abbreviations are used: Scen: scenario considered; AT/Cl: clean accuracy for the adversarially-pretrained model used as the classifier, when considered alone; C/Cl: clean accuracy for the Carso architecture; AT/AA: robust accuracy (by the means of AutoAttack) for the adversarially-pretrained model used as the classifier, when considered alone; C/randAA: robust accuracy for the Carso architecture, when attacked end-to-end by AutoAttack for randomised defences; Best AT/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial training alone (any model); Best P/AA: best robust accuracy result for the respective dataset (by the means of AutoAttack), obtained by adversarial purification (any model). Robust accuracies in round brackets are obtained using the Pgd+EoT (Lee & Kim, 2024) pipeline, developed for diffusion-based purifiers. The best clean and robust accuracies per dataset are shown in bold. The clean accuracies for the models referred to in the Best columns are shown in Table 19 (in Appendix E).
@@
Comparing the robust accuracy of the classifier model used in scenario (a) (Cui et al., 2023) with that resulting from the inclusion of the same model in the Carso architecture, we observe a $8.4\%$ increase. This is counterbalanced by a $5.6\%$ clean accuracy decrease. The same version of Carso further provides a +2.42+2.42 robustness increase w.r.t. the current best AT-trained model (Bartoldson et al., 2024) that employs a $\sim 4\times$ larger WideResNet-96-16 model.
In addition, our method provides a remarkable $9.72\%$ increase in robust accuracy w.r.t. to the best adversarial purification approach (Lin et al., 2024), a diffusion-based purifier. However, the comparison is not as straightforward. In fact, the original paper (Lin et al., 2024) reports a robust accuracy of $78.12\%$ using AutoAttack on the gradients obtained via the adjoint method (Nie et al., 2022). As noted in Lee & Kim (2024), such evaluation (which uses the version of AutoAttack that is unsuitable for stochastic defences) leads to a large overestimation of the robustness of diffusive purifiers. As suggested in Lee & Kim (2024), the authors of Lin et al. (2024) re-evaluate the robust accuracy according to a more suitable pipeline (Pgd+EoT, whose hyperparameters are shown in Table 14), obtaining a much lower robust accuracy of $66.41\%$. Consequently, we repeat the same evaluation for Carso and compare the worst-case robustness amongst the two. In line with typical AT methods, and unlike diffusive purification, the robustness of Carso assessed by means of randAA remains lower w.r.t. that achieved by Pgd+EoT.
@@
Moving to scenario (b), Carso achieves a robust accuracy increase of $27.47\%$ w.r.t. the classifier alone (Cui et al., 2023), balanced by a $5.79\%$ decrease in clean accuracy. Our approach also improves upon the robust accuracy of the best AT-trained model (Wang et al., 2023) (WideResNet-70-16) by $23.98\%$. In the absence of a reliable robustness evaluation by means of Pgd+EoT for the best purification-based method (Lin et al., 2024), we still obtain a $20.25\%$ increase in robust accuracy upon its (largely overestimated) AA result.
@@
In scenario (c), Carso improves upon the classifier alone (Wang et al., 2023) (which is also the best AT-based approach for TinyImageNet-200) by $22.26\%$. A significant clean accuracy toll is imposed by the relative complexity of the dataset, i.e. $8.87\%$. In this setting, we lack any additional purification-based methods.
@@
Although the architecture of Carso is algorithmically differentiable end-to-end – and the integrated diagnostics of the randAA routines identified no pitfalls during the assessment – we additionally guard against the eventual gradient obfuscation (Athalye et al., 2018a) induced by our method by repeating the evaluation at $\epsilon_{\infty}=0.95$, verifying that the resulting robust accuracy stays below random chance (Carlini et al., 2019). Results are shown in Table 3.
@@
Scenario (a) (b) (c) $\epsilon_{\infty}=0.95$ acc. <<0.047 <<0.010 $\approx$0.0